# Penrose limits and TsT for fibered $I$-branes

### v2

@@ 4.1.1 TsT transformation ($T\bar{T}$ deformation?) over $v,X$
The TsT transformation that we did in the last section, before doing the Penrose limit, was over the coordinates $t$ and $x$, and since we want to compare what are the effects of doing a TsT after the Penrose limit, we naturally need to deform in these same coordinates. However, the metric we obtained (4.3) is in terms of the more useful coordinates for the Penrose limit $u,v$, and since these coordinates are related to $t$ via $t=\frac{v}{2}-\frac{u}{2L^{2}}$, and for large $L$ the one that prevails is $v$, we do a TsT over $v$ and $X\propto x$.
After TsT transforming with parameter $\eta$, the metric (4.3) gets a correction of the order of $\eta^{2}$, and therefore the full metric can be written as $\tilde{ds}^{2}=ds^{2}+\eta^{2}\delta ds^{2}$, with $\delta ds^{2}$ being
@@
$B=\frac{\eta r_{+}^{2}}{2e_{A}e_{B}}(-e_{B}\Theta_{A}d\Phi_{A}+e_{A}\Theta_{B}d\Phi_{B}+e_{A}e_{B}du)\wedge dX.$
@@
$+dv\left(\frac{r_{+}\Theta_{B}}{e_{B}}\,d\Phi_{B}-\frac{r_{+}\Theta_{A}}{e_{A}}\,d\Phi_{A}\right)+\eta^{2}\delta ds^{2}.$
@@
$ds^{2}$ $=$ $\frac{1}{4}\Bigg(-\frac{r_{+}\left(-8+\eta^{2}r^{2}\Theta_{A}^{2}\right)d\Phi_{A}^{2}}{e_{A}^{2}}-\frac{r_{+}\left(-8+\eta^{2}r^{2}\Theta_{B}^{2}\right)d\Phi_{B}^{2}}{e_{B}^{2}}+16r_{+}\,d\Theta^{2}$ (4.16)
$+\frac{8r_{+}\,d\Theta_{A}^{2}}{e_{A}^{2}}+\frac{8r_{+}\,d\Theta_{B}^{2}}{e_{B}^{2}}-\eta^{2}r^{3}\,du^{2}-\frac{2r_{+}\Theta_{B}\,d\Phi_{B}\left(\eta^{2}r_{+}^{2}du-2dv\right)}{e_{B}}+4r_{+}\,du\,dv$
$+\frac{2r_{+}\Theta_{A}\,d\Phi_{A}\left(\eta^{2}r_{+}^{2}\Theta_{B}\,d\Phi_{B}+e_{B}\eta^{2}r_{+}^{2}du-2e_{B}dv\right)}{e_{A}e_{B}}+4r_{+}\,dX^{2}$
@@
$\varphi\to\frac{4\,\varphi}{r_{+}f_{s}^{\prime}},\Theta_{A}\to\frac{e_{A}\,\Theta_{A}}{\sqrt{2r_{+}}},\Theta_{B}\to\frac{e_{B}\,\theta_{B}}{\sqrt{2r_{+}}},u\to\frac{2(v-u)}{\eta r_{+}^{3/2}},v\to-\eta\sqrt{r}\,u\;,$
@@
$+\left(1-\frac{1}{8}\eta^{2}r_{+}^{2}\Theta_{A}^{2}\right)\,d\Phi_{A}^{2}+\left(1-\frac{1}{8}\eta^{2}r_{+}^{2}\Theta_{B}^{2}\right)\,d\Phi_{B}^{2}$
$+\frac{1}{4}\eta^{2}r_{+}^{2}\Theta_{A}\Theta_{B}\,d\Phi_{A}\,d\Phi_{B}+d\Theta_{A}^{2}+d\Theta_{B}^{2}$
$-dv^{2}+du^{2}+\frac{\eta r_{+}\Theta_{A}\,d\Phi_{A}\,dv}{\sqrt{2}}-\frac{\eta r_{+}\Theta_{B}\,d\Phi_{B}\,dv}{\sqrt{2}}.$
@@ 4.1.2 TsT transformation over $v,\varphi$ and $v,u$
$M=\left(1-\eta(e_{v\varphi}-e_{\varphi v}\right)+\eta^{2}\det\left(\begin{pmatrix}e_{vv}&e_{v\varphi}\\ e_{\varphi v}&e_{\varphi\varphi}\end{pmatrix}\right)^{-1}\;,$ (4.22)
@@
$\quad-\frac{\eta^{2}\rho^{2}r_{+}^{3}f_{s}^{\prime}(r_{+})}{Y}du^{2}+\frac{4\rho^{2}r_{+}f_{s}^{\prime}(r_{+})}{Y}d\varphi^{2}$
$\quad+\frac{2r_{+}Y-\eta^{2}\rho^{2}r_{+}^{3}\theta_{A}^{2}f_{s}^{\prime}(r_{+})}{e_{a}^{2}Y}d\phi_{A}^{2}+\frac{2r_{+}Y-\eta^{2}\rho^{2}r_{+}^{3}\theta_{B}^{2}f_{s}^{\prime}(r_{+})}{e_{b}^{2}Y}d\phi_{B}^{2}$
@@
$\quad+du\left(\frac{2\eta^{2}\rho^{2}r_{+}^{3}\Theta_{A}f_{s}^{\prime}(r_{+})}{e_{a}Y}d\Phi_{A}-\frac{2\eta^{2}\rho^{2}r_{+}^{3}\Theta_{B}f_{s}^{\prime}(r_{+})}{e_{b}Y}d\Phi_{B}\right)$
$\quad+\frac{2\eta^{2}\rho^{2}r_{+}^{3}\Theta_{A}\Theta_{B}f_{s}^{\prime}(r_{+})}{e_{a}e_{b}Y}d\Phi_{A}d\Phi_{B}+\frac{4r_{+}}{Y}du\,dv\;,$
where $Y=4-\eta^{2}\rho^{4}(Q_{A}-Q_{B})^{2}r_{+}^{2}\zeta^{\prime}(r_{+})^{2}$, and there are also nonzero B and $\phi(\rho)$ fields present. It is important to emphasize that the dilaton $\phi$ becomes radial dependent, implying an energy scale dependence in the field theory. Again, the $C_{2}$ field remains the same, and we have a $C_{4}$ just like in the previous cases. The form of the metric suggest again that the TsT deformation is dual to an irrelevant operator, because as $\rho\rightarrow 0$, it reduces again to the undeformed pp wave metric. The general structure of the solution does not easily suggest a field theory interpretation, since it is a complex combination of coordinates with a not so obvious meaning from the dual side. However, we can try to evaluate the deformation of the special case $Q_{A}=Q_{B}$ of the pp wave, where (if also $e_{A}=e_{B}$) $24$ supersymmetries were preserved. Under this condition, the metric reduces to the simple form
@@
$-\frac{\eta^{2}\rho^{2}r_{+}^{3}f_{s}^{\prime}(r_{+})}{4e_{A}^{2}e_{B}^{2}}\left(e_{A}e_{B}\,du+e_{A}\Theta_{B}\,d\Phi_{B}-e_{B}\Theta_{A}\,d\Phi_{A}\right)^{2}\;,$
@@
$B=\frac{\eta\rho^{2}r_{+}^{2}f_{s}^{\prime}(r_{+})}{2e_{A}e_{B}}\left(e_{B}\Theta_{A}d\Phi_{A}\wedge d\varphi-e_{A}\Theta_{B}d\Phi_{B}\wedge d\varphi-e_{A}e_{B}du\wedge d\varphi\right)\;,$ (4.30)
@@
$ds^{2}$ $=$ $-\frac{2}{\eta^{2}-1}dudV+\frac{1}{\eta^{2}-1}dV^{2}\left(\frac{e_{B}^{2}}{16}r_{+}|\omega_{1}|^{2}+\frac{e_{A}^{2}}{16}r_{+}|\omega_{2}|^{2}+\frac{(Q_{B}-Q_{A})^{2}\zeta^{\prime}(r_{+})^{2}}{16}|\omega_{3}|^{2}\right)$ (4.34)
@@
$B_{2}$ $=$ $\frac{\eta}{\eta^{2}-1}du\wedge dv$ (4.35)
$\Phi$ $=$ $\log r_{+}-\frac{1}{2}\log(1-\eta^{2}).$ (4.36)
@@
$=$ $\frac{\eta V}{2(1-\eta^{2})r_{+}^{2}f_{s}^{\prime}(r_{+})}(-e_{B}f_{s}^{\prime}(r_{+})d\Phi_{B}\wedge d\Theta_{B}+e_{A}f_{s}^{\prime}(r_{+})d\Phi_{A}\wedge d\Theta_{A}$ (4.39)
@@ 4.1.3 General cases for TsT transformation
    $\delta ds^{2}=-\frac{\eta^{2}r_{+}^{3}}{4e_{A}^{2}e_{B}^{2}}(-e_{B}\Theta_{A}d\Phi_{A}+e_{A}\Theta_{B}d\Phi_{B}+e_{A}e_{B}du)^{2}.$ (4.40)
@@
    $\delta ds^{2}=-\frac{\eta^{2}\rho^{2}r_{+}^{3}f_{s}^{\prime}(r_{+})}{4e_{A}^{2}e_{B}^{2}}(-e_{B}\Theta_{A}d\Phi_{A}+e_{A}\Theta_{B}d\Phi_{B}+e_{A}e_{B}du)^{2}.$ (4.41)
@@
    $\delta ds^{2}=-\frac{\eta^{2}r_{+}^{3}}{e_{A}^{2}e_{B}^{2}}(-e_{B}\Theta_{A}d\Phi_{A}+e_{A}\Theta_{B}d\Phi_{B}+e_{A}e_{B}du)^{2}.$ (4.42)
@@
    $B=\frac{\eta r_{+}^{2}}{2e_{A}e_{B}}(-e_{B}\Theta_{A}d\Phi_{A}+e_{A}\Theta_{B}d\Phi_{B}+e_{A}e_{B}du)\wedge dX.$ (4.44)
@@
    $B=\frac{\eta\rho^{2}r_{+}^{2}}{2e_{A}e_{B}}f_{s}^{\prime}(r_{+})(-e_{B}\Theta_{A}d\Phi_{A}+e_{A}\Theta_{B}d\Phi_{B}+e_{A}e_{B}du)\wedge d\varphi.$ (4.45)
@@
    $B=\frac{\eta r_{+}^{2}}{2e_{A}e_{B}}(-e_{B}\Theta_{A}d\Phi_{A}+e_{A}\Theta_{B}d\Phi_{B}+e_{A}e_{B}du)\wedge d\Theta.$ (4.46)
@@
Moving on, there is another interesting relation among the TsT deformations, when the pair of coordinates involves the $\varphi$ one. As we saw before, the TsT over $(v,\varphi)$ generates a deformation in the gauge side analogous to what would be the deformation done by an irrelevant operator. We now show that this is not a coincidence: any TsT over $\varphi$ generates dual gauge theories deformed by an irrelevant operator (since we already studied $(v,\varphi)$, we will ignore it here),
@@
Besides that, we see two interesting facts regarding these deformations: while the $(\Theta,\varphi)$ and $(X,\varphi)$ cases vanish for $\rho\rightarrow 0$, and the dilaton goes to a constant, for $\rho\rightarrow\infty$ the B field goes to a constant (and therefore, can be gauged away; though one could still have noncommutativity, in principle), and the dilaton goes to $\phi\rightarrow-\log\rho$, which means that the coupling constant $g_{s}=e^{\phi}\sim\frac{1}{\rho}\rightarrow 0$, meaning that there is no noncommutativity, and therefore the deformed gauge side is free, and perturbative at high energies, with operators interacting via the usual product. This is a remarkable result, since this suggests that we have an asymptotically free gauge theory, at least in some large R-charge sector.

### v4

@@ 4.1.1 TsT transformation ($T\bar{T}$ deformation?) over $v,X$
The TsT transformation that we did in the last section, before doing the Penrose limit, was over the coordinates $t$ and $x$, and since we want to compare what are the effects of doing a TsT after the Penrose limit, we naturally need to deform in these same coordinates. However, the metric we obtained (4.3) is in terms of the more useful coordinates for the Penrose limit $u,v$, and since these coordinates are related to $t$ via $t=\frac{v}{2}-\frac{u}{2L^{2}}$, and for large $L$ the one that prevails is $v$, we do a TsT over $v$ and $X\propto x$. However, since $v$ is now the timelike coordinate, we again Wick rotate it, do the TsT and then Wick rotate back.
After TsT transforming with parameter $\eta$, the metric (4.3) gets a correction of the order of $\eta^{2}$, and therefore the full metric can be written as $\tilde{ds}^{2}=ds^{2}-\eta^{2}\delta ds^{2}$, with $\delta ds^{2}$ being
@@
$B=-\frac{\eta r_{+}^{2}}{2e_{A}e_{B}}(-e_{B}\Theta_{A}d\Phi_{A}+e_{A}\Theta_{B}d\Phi_{B}+e_{A}e_{B}du)\wedge dX.$
@@
$+dv\left(\frac{r_{+}\Theta_{B}}{e_{B}}\,d\Phi_{B}-\frac{r_{+}\Theta_{A}}{e_{A}}\,d\Phi_{A}\right)-\eta^{2}\delta ds^{2}.$
@@
$ds^{2}$ $=$ $\frac{1}{4}\Bigg(-\frac{r_{+}\left(-8-\eta^{2}r^{2}\Theta_{A}^{2}\right)d\Phi_{A}^{2}}{e_{A}^{2}}-\frac{r_{+}\left(-8-\eta^{2}r^{2}\Theta_{B}^{2}\right)d\Phi_{B}^{2}}{e_{B}^{2}}+16r_{+}\,d\Theta^{2}$ (4.16)
$+\frac{8r_{+}\,d\Theta_{A}^{2}}{e_{A}^{2}}+\frac{8r_{+}\,d\Theta_{B}^{2}}{e_{B}^{2}}+\eta^{2}r^{3}\,du^{2}-\frac{2r_{+}\Theta_{B}\,d\Phi_{B}\left(+\eta^{2}r_{+}^{2}du-2dv\right)}{e_{B}}+4r_{+}\,du\,dv$
$+\frac{2r_{+}\Theta_{A}\,d\Phi_{A}\left(-\eta^{2}r_{+}^{2}\Theta_{B}\,d\Phi_{B}-e_{B}\eta^{2}r_{+}^{2}du-2e_{B}dv\right)}{e_{A}e_{B}}+4r_{+}\,dX^{2}$
@@
$\varphi\to\frac{4\,\varphi}{r_{+}f_{s}^{\prime}},\Theta_{A}\to\frac{e_{A}\,\Theta_{A}}{\sqrt{2r_{+}}},\Theta_{B}\to\frac{e_{B}\,\theta_{B}}{\sqrt{2r_{+}}},u\to\frac{2(v-u)}{\eta r_{+}^{3/2}},v\to+\eta\sqrt{r}\,u\;,$
@@
$+\left(1+\frac{1}{8}\eta^{2}r_{+}^{2}\Theta_{A}^{2}\right)\,d\Phi_{A}^{2}+\left(1+\frac{1}{8}\eta^{2}r_{+}^{2}\Theta_{B}^{2}\right)\,d\Phi_{B}^{2}$
$-\frac{1}{4}\eta^{2}r_{+}^{2}\Theta_{A}\Theta_{B}\,d\Phi_{A}\,d\Phi_{B}+d\Theta_{A}^{2}+d\Theta_{B}^{2}$
$-dv^{2}+du^{2}-\frac{\eta r_{+}\Theta_{A}\,d\Phi_{A}\,dv}{\sqrt{2}}+\frac{\eta r_{+}\Theta_{B}\,d\Phi_{B}\,dv}{\sqrt{2}}.$
@@ 4.1.2 TsT transformation over $v,\varphi$ and $v,u$
$M=\left(1+\eta(e_{v\varphi}-e_{\varphi v}\right)-\eta^{2}\det\left(\begin{pmatrix}e_{vv}&e_{v\varphi}\\ e_{\varphi v}&e_{\varphi\varphi}\end{pmatrix}\right)^{-1}\;,$ (4.22)
@@
$\quad+\frac{\eta^{2}\rho^{2}r_{+}^{3}f_{s}^{\prime}(r_{+})}{Y}du^{2}+\frac{4\rho^{2}r_{+}f_{s}^{\prime}(r_{+})}{Y}d\varphi^{2}$
$\quad+\frac{2r_{+}Y+\eta^{2}\rho^{2}r_{+}^{3}\theta_{A}^{2}f_{s}^{\prime}(r_{+})}{e_{a}^{2}Y}d\phi_{A}^{2}+\frac{2r_{+}Y+\eta^{2}\rho^{2}r_{+}^{3}\theta_{B}^{2}f_{s}^{\prime}(r_{+})}{e_{b}^{2}Y}d\phi_{B}^{2}$
@@
$\quad+du\left(\frac{-2\eta^{2}\rho^{2}r_{+}^{3}\Theta_{A}f_{s}^{\prime}(r_{+})}{e_{a}Y}d\Phi_{A}-\frac{-2\eta^{2}\rho^{2}r_{+}^{3}\Theta_{B}f_{s}^{\prime}(r_{+})}{e_{b}Y}d\Phi_{B}\right)$
$\quad-\frac{2\eta^{2}\rho^{2}r_{+}^{3}\Theta_{A}\Theta_{B}f_{s}^{\prime}(r_{+})}{e_{a}e_{b}Y}d\Phi_{A}d\Phi_{B}+\frac{4r_{+}}{Y}du\,dv\;,$
where $Y=4+\eta^{2}\rho^{4}(Q_{A}-Q_{B})^{2}r_{+}^{2}\zeta^{\prime}(r_{+})^{2}$, and there are also nonzero B and $\phi(\rho)$ fields present. It is important to emphasize that the dilaton $\phi$ becomes radial dependent, implying an energy scale dependence in the field theory. Again, the $C_{2}$ field remains the same, and we have a $C_{4}$ just like in the previous cases. The form of the metric suggest again that the TsT deformation is dual to an irrelevant operator, because as $\rho\rightarrow 0$, it reduces again to the undeformed pp wave metric. The general structure of the solution does not easily suggest a field theory interpretation, since it is a complex combination of coordinates with a not so obvious meaning from the dual side. However, we can try to evaluate the deformation of the special case $Q_{A}=Q_{B}$ of the pp wave, where (if also $e_{A}=e_{B}$) $24$ supersymmetries were preserved. Under this condition, the metric reduces to the simple form
@@
$+\frac{\eta^{2}\rho^{2}r_{+}^{3}f_{s}^{\prime}(r_{+})}{4e_{A}^{2}e_{B}^{2}}\left(e_{A}e_{B}\,du+e_{A}\Theta_{B}\,d\Phi_{B}-e_{B}\Theta_{A}\,d\Phi_{A}\right)^{2}\;,$
@@
$B=-\frac{\eta\rho^{2}r_{+}^{2}f_{s}^{\prime}(r_{+})}{2e_{A}e_{B}}\left(e_{B}\Theta_{A}d\Phi_{A}\wedge d\varphi-e_{A}\Theta_{B}d\Phi_{B}\wedge d\varphi-e_{A}e_{B}du\wedge d\varphi\right)\;,$ (4.30)
@@
$ds^{2}$ $=$ $+\frac{2}{\eta^{2}+1}dudV-\frac{1}{\eta^{2}+1}dV^{2}\left(\frac{e_{B}^{2}}{16}r_{+}|\omega_{1}|^{2}+\frac{e_{A}^{2}}{16}r_{+}|\omega_{2}|^{2}+\frac{(Q_{B}-Q_{A})^{2}\zeta^{\prime}(r_{+})^{2}}{16}|\omega_{3}|^{2}\right)$ (4.34)
@@
$B_{2}$ $=$ $\frac{\eta}{\eta^{2}+1}du\wedge dv$ (4.35)
$\Phi$ $=$ $\log r_{+}-\frac{1}{2}\log(1+\eta^{2}).$ (4.36)
@@
$=$ $\frac{-\eta V}{2(1+\eta^{2})r_{+}^{2}f_{s}^{\prime}(r_{+})}(-e_{B}f_{s}^{\prime}(r_{+})d\Phi_{B}\wedge d\Theta_{B}+e_{A}f_{s}^{\prime}(r_{+})d\Phi_{A}\wedge d\Theta_{A}$ (4.39)
@@ 4.1.3 General cases for TsT transformation
    $\delta ds^{2}=+\frac{\eta^{2}r_{+}^{3}}{4e_{A}^{2}e_{B}^{2}}(-e_{B}\Theta_{A}d\Phi_{A}+e_{A}\Theta_{B}d\Phi_{B}+e_{A}e_{B}du)^{2}.$ (4.40)
@@
    $\delta ds^{2}=+\frac{\eta^{2}\rho^{2}r_{+}^{3}f_{s}^{\prime}(r_{+})}{4e_{A}^{2}e_{B}^{2}}(-e_{B}\Theta_{A}d\Phi_{A}+e_{A}\Theta_{B}d\Phi_{B}+e_{A}e_{B}du)^{2}.$ (4.41)
@@
    $\delta ds^{2}=+\frac{\eta^{2}r_{+}^{3}}{e_{A}^{2}e_{B}^{2}}(-e_{B}\Theta_{A}d\Phi_{A}+e_{A}\Theta_{B}d\Phi_{B}+e_{A}e_{B}du)^{2}.$ (4.42)
@@
    $B=-\frac{\eta r_{+}^{2}}{2e_{A}e_{B}}(-e_{B}\Theta_{A}d\Phi_{A}+e_{A}\Theta_{B}d\Phi_{B}+e_{A}e_{B}du)\wedge dX.$ (4.44)
@@
    $B=-\frac{\eta\rho^{2}r_{+}^{2}}{2e_{A}e_{B}}f_{s}^{\prime}(r_{+})(-e_{B}\Theta_{A}d\Phi_{A}+e_{A}\Theta_{B}d\Phi_{B}+e_{A}e_{B}du)\wedge d\varphi.$ (4.45)
@@
    $B=-\frac{\eta r_{+}^{2}}{2e_{A}e_{B}}(-e_{B}\Theta_{A}d\Phi_{A}+e_{A}\Theta_{B}d\Phi_{B}+e_{A}e_{B}du)\wedge d\Theta.$ (4.46)
@@
Moving on, there is another interesting relation among the TsT deformations, when the pair of coordinates involves the $\varphi$ one. As we saw before, the TsT over $(v,\varphi)$ generates a deformation in the gauge side analogous to what would be the deformation done by an irrelevant operator. We now show that this is not a coincidence: any TsT over $\varphi$ generates dual gauge theories deformed by an irrelevant operator (since we already studied $(v,\varphi)$, we will ignore it here),  [^6] Note, however, that since in these cases there was no temporal coordinate ($v$) involved, there was no need to make any Wick rotation. Indeed, there are no apparent singularities appearing.
@@
Besides that, we see two interesting facts regarding these deformations: while the $(\Theta,\varphi)$ and $(X,\varphi)$ cases vanish for $\rho\rightarrow 0$, and the dilaton goes to a constant, for $\rho\rightarrow\infty$ the B field goes to a constant (and therefore, can be gauged away; though one could still have noncommutativity, in principle), and the dilaton goes to $\phi\rightarrow-\log\rho$, which means that the coupling constant $g_{s}=e^{\phi}\sim\frac{1}{\rho}\rightarrow 0$, meaning that there is no noncommutativity, and therefore the deformed gauge side is free, and perturbative at high energies, with operators interacting via the usual product. This is a remarkable result, since this suggests that we have an asymptotically free gauge theory, at least in some large R-charge sector. It is, however, true that the dilaton alone does not define the gauge coupling in these cases, one has to consider a probe DBI brane action in the background for that, as in [1]. We do this in Appendix A.4, where we find that the behaviour of the dilaton is qualitatively the same as of the gauge coupling, at least in the case of the TsT transformed I-brane background, so the conclusion stands.